Ruth is x years old. In 12 years time, Ruth will be three times as old as she was 6 years ago. Find Ruth's age now.
Answer by addingup(3677)

x+12 = 3(x-6)
x+12 = 3x-18 Subtract 3x and 12 from both sides
-2x = -30 divide both sides by -2, and remember -/- = +
x = 15 Ruth is 15
------------------------------------
Check:
15+12 = 3(15-6)
15+12 = 45-18
27 = 27 Correct.
:
John